OpenStudy (anonymous):

Jeff is baking a cake. The recipe says that he has to mix 32 grams of vanilla powder to the flour. Jeff knows that 1 cup of that particular vanilla powder has a mass of 128 grams. He added two over three of a cup of vanilla powder to the flour. Should Jeff add more vanilla powder to make the exact recipe or did he go over and by what amount?

OpenStudy (anonymous):

He already added: \[\frac{ 2 }{ 3 }of.a.cup\]do you know how to find the amount of gram that he already added?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

this is how to look at it: if one cup is 128g how much is \[\frac{ 3 }{ 3 }of.a.cup\]?

OpenStudy (anonymous):

3/3=1 soo would it be 128?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to know the amount he already added 2/3 * 128g

OpenStudy (anonymous):

85.3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

he should actually add 32g , so he went over by how much?

OpenStudy (anonymous):

53.3

OpenStudy (anonymous):

and that's all answers to the question

OpenStudy (anonymous):

the only options i have are 5/12 or 1/4 ..?

OpenStudy (anonymous):

ok they don't want us to use the calculator for everything the amount he should add compared with a cup is 32g / 128 g can you reduce that fraction ?

OpenStudy (anonymous):

its 1/4.! Gracias!! : )

OpenStudy (anonymous):

he should have added 1/4

OpenStudy (anonymous):

however he already added 2/3 cups so since he was supposed to add: 1/4 = 3/12 but he added 2/3 = 8/12 he's 5/12 over the recipe

OpenStudy (anonymous):

ohhhhh, i see.